OpenStudy (anonymous):

Example: Convert 25°C to °F. Formula: °F = (9/5) (°C) + 32  °C is given so I am solving for °F. °F = (9/5) (25) + 32 =(1.8*25)+ 32 =45+32 =77 Answer= 77°F is equal to 25°C A. Convert 35°C to °F B. Convert 80°F to °C C. Convert 15°C to °F D. Convert 25°F to °C

OpenStudy (anonymous):

@Mehek14 @Mashy @jhonyy9

jhonyy9 (jhonyy9):

so there you have an example what you can using easy - try use this way to solve the next transformations - ok. ?

OpenStudy (anonymous):

okay :)

jhonyy9 (jhonyy9):

Answer= 77°F is equal to 25°C A. Convert 35°C to °F B. Convert 80°F to °C C. Convert 15°C to °F D. Convert 25°F to °C or than you like you can using the formula of 3 simple what say than 77°F ----- 25°C x°F ----- 35°C --------------- cross multiplie and solve for x - how many will get for x ?

jhonyy9 (jhonyy9):

cross multiplie and will get 25x=35*77 yes ? from these how you calculi the x ?

jhonyy9 (jhonyy9):

divide both sides by 25 and will get x = 35*77/25 simplifie 35 and 25 by 5 and will get x = 7*77/5 = ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

this really confusing

OpenStudy (anonymous):

You have a formula for calculating temperatures in Farenheit, given temperatures in Celsius: F = (9/5)*C + 32 F is solved for; your input is temperature in Celsius and your output is temperature in Farenheit. To go the other way, from Celsius to Farenheit, you need to algebraically manipulate the formula and solve for Celsius. Here's the first step: F - 32 = (9/5)*C I subtracted 32 from both sides of the equation, so now there is a negative 32 on the left-hand side. Do you see what you can do next to solve for C? Then you will have temperature in Celsius as your output, and degrees Farenheit as your input.

OpenStudy (anonymous):

@etcyl

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

What is confusing you?